OpenStudy (anonymous):

The local community park is designing a pool using the Golden Ratio. If the pool is 60 feet wide, about how long should it be? Answer 60 feet 100 feet 120 feet 150 feet

OpenStudy (anonymous):

120 feet?

OpenStudy (anonymous):

it should be 100 because the golden ratio is 1.61803399

Directrix (directrix):

The golden ratio "phi" has approximate value 1.61803399.

Directrix (directrix):

60 is to the length L as 1 is to phi 60/L = 1/1.61803399 L = 60 (1.61803399 ) = 97.082 approx

Directrix (directrix):

100/60 = 1.6 with a repeating 6 so I'd go with 100 as the answer.